# On the attribution of weather events to climate change using a fit to extreme value distributions

### v2

@@ 2.6 Data
Following CC2021, we rely on two sets of observations: (1) daily maximum temperature extremes for each year from station data in Verkhoyansk as discussed in CC2021; and (2) 4-year smoothed GMST anomalies from GISTEMP Lenssen et al. (2019). CC2021 also performed a GEV analysis on temperature anomalies over a region in Siberia, which we do not focus on here in order to study specifically the extreme values detected at the Verkhoyansk station. At the time of the CC2021 study, the daily maximum temperature extreme in 2020 at the Verkhoyansk station was recorded as 38C. This temperature extreme currently appears as a missing data point on the National Climatic Data Center website (https://www.ncei.noaa.gov/cdo-web/search, station RSM00024266, VERHOJANSK, data downloaded Aug 2023) but we still use this unconfirmed value in our analysis. We also study GMST anomalies and extreme temperatures over the Siberian region defined in CC2021 within the CESM Large Ensemble Kay et al. (2015) over the period 1926–2019. The CESM Large Ensemble contains 42 simulations of the climate over the period 1920–2100 under historical and RCP 8.5 external forcing, of which we use the 35 that were run on the NCAR supercomputer. We use model output only up to the year 2020, and as a result, there should be little sensitivity to the emissions scenario used (DAMIP, Gillett et al., 2016, uses RCP 4.5), as the emissions scenarios begin to significantly diverge only later. These simulations differ only in small perturbations to the initial conditions, which lead to different temporal sequences of internal variability. Similarly, we study analogous variables within a 1,800-year CESM preindustrial simulation, yielding 18 segments of 94 years for consistency with the length of the observed record.
Australian Bushfire.
Following OK2021, we again use two sets of observations for the years 1920–2019: (1) annual (July–June) maxima of 7-day moving average daily-maximum surface temperature data from the Australian Water Availability Project (AWAP) over a region in Australia defined in OK2021; and (2) 4-year smoothed GMST anomalies from GISTEMP. We again apply the same analysis to the CESM Large Ensemble and to the long CESM preindustrial run, in this case using 18 segments of 98 years, for consistency with the observations.
@@
Following HW2022, we used two datasets. (1) 2-year means of precipitation from ERA5 Hersbach et al. (2020) over the region of Madagascar that was defined in HW2022. HW2022 explained that they used a “24-month running mean rainfall data from July to June.” Our approach is to calculate total annual precipitation and then calculate a running 2-year average of this annual data resulting in a new smoothed annual time series. For 1951–2020 the bottom 20% precipitation 2-year periods are selected for the analysis, corresponding to the following 14 two-year periods: 1956–1957, 1957–1958, 1958–1959, 1959–1960, 1962–1963, 1990–1991, 1991–1992, 1992–1993, 2008–2009, 2009–2010, 2015–2016, 2016–2017, 2017–2018, 2019–2020. (2) 4-year smoothed GMST anomalies from GISTEMP. We note that the analysis following HW2022 relies on a fit to only 14 data points. We discuss the issues that are involved in fitting extreme distribution functions to such a small number of data points below. We also study the equivalent variables within the CESM Large Ensemble over the period 1950–2019, consistent with the period of observations. We again apply the same analysis to the CESM Large Ensemble and to the long CESM preindustrial run, in this case using 18 segments of 70 years.
WWA uses two variants of their analysis approach, one in which the extreme data point being attributed is incorporated into the estimate of the PDFs and one in which it is not Philip et al. (2020). We chose to use the second approach. Inclusion of the extreme being analyzed in the calculation will tend to make the parametric dependence on GMST larger, and tend to make it easier to reject the null of no change in favor of the alternate hypothesis. In the present case, we are still unable to reject the null hypothesis, such that the exclusion of the event in question will not change the result.
We note that in each of these three case studies, WWA leverages many datasets, including ERA5 reanalysis, station observations, and the multimodel CMIP5 ensemble. We use only a subset of the data they study and also the CESM Large Ensemble data for self-consistency in comparing the three examples that we analyze here.
@@ 3 Results
Our analysis of the attribution of extreme events based on a fit to extreme value distributions is presented as follows. First, in subsection 33.1, we assess the null hypothesis that a lack of influence of anthropogenic climate change should result in no GMST dependence of the distribution parameters or in $\alpha$ being indistinguishable from zero. We do so by analyzing a 1,800-year CESM preindustrial simulation. The preindustrial simulation uses fixed greenhouse gas concentrations, volcanic activity, anthropogenic aerosols, land use, and solar forcing all set to their respective levels in the year 1850. Most importantly in the context of this study is the lack of a time-varying anthropogenic greenhouse gas forcing in this simulation. In the following three sections (33.2, 33.3, 33.4) we then consider each of the three test cases and first assess the need for a GMST dependence and also for using a GEV or GPD fit to the observations as opposed to that of a normal or log-normal distribution. We then evaluate whether there is a statistically significant GMST dependence using the deviance statistic. Finally, we estimate the uncertainty level of the parameter $\alpha$ used for the attribution and evaluate the GEV or GPD fit to the CESM Large Ensemble in order to deduce the amount of data needed to adequately constrain GMST dependence.
3.1 Extremes analysis of a pre-industrial simulation
@@
WWA examines the dependence of the distribution parameters of extreme events on the (4-year running mean) GMST and, if $\alpha$ is significantly greater than zero, to infer that anthropogenic climate change affects these extreme events. Furthermore, the parameter $\alpha$ representing the dependency on GMST is then used to quantify how much a given extreme event was made more likely as a consequence of climate change. We show here that it is possible, however, for internal climate variability, such as ENSO, to produce covariance between GMST and extremes even without anthropogenic influences. This is demonstrated most clearly in this section by using a pre-industrial climate model run, and we further elaborate using the CESM Large Ensemble as well. More importantly, for attribution purposes, this would suggest that the value of $\alpha$ calculated in the presence of a GHG increase can be affected by natural variability, biasing the attribution results.
@@
To examine this potential role of internal variability, we start with the case of the hot weather associated with the Australian Bushfires and perform a GEV analysis using a 1,800-year CESM preindustrial run. This run is done without time-varying changes to anthropogenic CO2, other anthropogenic greenhouse emissions, volcanic (natural) emissions, land use changes, or aerosols. These are all held constant at year 1850 values. We use 18 data segments of length 94 years, equal to the length of the observed record and analogous to having 18 simulations of historical climate with no changing anthropogenic emissions. The distribution of $\alpha$ values for the GEV preindustrial analysis are shown by the green bars in Fig. 1a. Despite no increase to anthropogenic forcing being applied to the preindustrial run, we find a large and positive $\alpha$ calculated from essentially all model output segments centered around a median value of 4.9 C per C. The large $\alpha$ value means that for every 1 C GMST warming, extreme summer temperatures increase by 4.9 C. Note that, following WWA, we are using four-year-running average GMST values for this computation.
@@
It is difficult to imagine a physical causal mechanism that would lead to such an enormous amplification of the GMST signal in Australian heatwaves. Instead, internal variability, for example ENSO, could increase both extreme temperatures over Australia and, to a much smaller degree, affect GMST, creating this seemingly strong dependence between extreme events and GMST. It is known that several modes of internal variability affect extreme temperatures in Australia, including ENSO, the Indian Ocean Dipole, and the Southern Annular Mode Hendon et al. (2007); Ummenhofer et al. (2009); King et al. (2020). The covariance between this extreme and GMST found in the CESM Large Ensemble analysis thus must result from another factor (internal climate variability, such as ENSO) affecting both Australian temperatures and global mean temperatures. This highlights a potential major difficulty with analyses based on a fit of distribution parameters to GMST. When analyzing the observed record in section 3.3 below, we find $\alpha\approx 2\;^{\circ}$ per C. This suggests a two-degree warming of extreme events for a one-degree warming of the GMST. Given the results here, it is more likely, however, that this estimate for $\alpha$ is influenced by natural variability as demonstrated by analyzing the preindustrial run with a time-invariant anthropogenic forcing.
@@
The same analysis for Siberian heatwaves (green bars in Fig. 1b) shows that the value for $\alpha$ deduced from a record of a similar length to that of the observations’ ranges between 1 to 4, although there is again no time-varying anthropogenic forcing in the corresponding model run. This large and variable response shows that a non-zero $\alpha$ does not necessarily imply an effect of anthropogenic climate change. The base hypothesis that a non-zero GMST dependence implies an anthropogenic signal is contradicted by the above analysis.
@@
For the Madagascar Drought example (Fig. 1c), we find a large spread in both the preindustrial run as well as the CESM Large Ensemble, centered around zero in both cases. The fact that the $\alpha$ distribution from the preindustrial run is centered around zero suggests that a null hypothesis of $\alpha=0$ could be appropriate in this case. That is, low extremes of precipitation show no significant increase or decrease related to GMST in either the forced or unforced runs. We further consider this case below using observations.
The results for Australian extremes (Figs. 1a) show that the distribution of $\alpha$ in increasing GHG forced simulations (blue bars) is closer to zero than in the unforced simulations (green bars). This demonstrates that $\alpha$ is affected by both GMST and by the weak projection of regional natural variability on the GMST. In the first case, a sensible null hypothesis is that the PDF of heat extremes is shifted by the warming in GMST as may be better represented by $\alpha=1$, while we saw that the second case can lead to large $\alpha$ values. In simulations with time-varying anthropogenic forcing, there is no way to tell what the effect of natural variability on $\alpha$ is using this approach, so the magnitude of $\alpha$ appears a poor indicator of anthropogenic influence.
@@
WWA recognizes that $\alpha$ may be strongly affected by natural variability and uses the 4-year running mean GMST to filter out the effects of natural variability. The analysis in this paper, however, shows that this approach is insufficient and that such variability can still lead to a large $\alpha$ even with smoothing. We conclude that in a more typical scenario analyzed using the WWA methodology and applied to observations under varying GHG, the value of $\alpha$ may still be affected by natural variability and, therefore, may bias the estimates of the effect of anthropogenic climate change on the return time of extreme events. One could justify the WWA use of the 4-year averaged GMST as being a proxy for known greenhouse gas concentrations. If CO2 were indeed used, the analysis of the preindustrial run, would, of course, lead to $\alpha=0$, eliminating this problem. But WWA studies use GMST and we follow that methodology here.
@@
The result that natural variability may strongly affect $\alpha$ when the 4-year average GMST is used to link extreme events to anthropogenic climate change suggests that it might be useful to reevaluate some of the WWA results based on the fit to extreme distributions in at least the three studies that we review. It is possible, of course, that a strong anthropogenic signal in a few decades could significantly affect extreme events and overwhelm the effects of natural variability in cases such as the Siberian Heatwave. The attribution of extreme temperatures would be more difficult in cases like the Australian example, where the effects of natural variability are large, as indicated by the large $\alpha$ detected in the CESM preindustrial simulation. In such cases, it would be important to correctly identify all relevant covariates (such as ENSO in the Australian case) and note that misleading results for $\alpha$ may be obtained if one relies only on GMST. It is, therefore, necessary to study when and how such an emergence of the ACC signal out of natural variability occurs for this attribution analysis to be used with confidence. We next proceed to consider other uncertainties involved in the attribution analysis based on a fit of extreme value distributions.
@@
Figure 1: Effect of natural climate variability on $\alpha$ distribution. PDFs of $\alpha$ for (a) The Australian 7-day moving mean of the maximum temperature related to the analysis of the 2019–2020 bushfires. (b) The Siberian maximum daily temperatures that were used in the analysis of the 2020 heatwave. (c) The Madagascar 2-year precipitation used in the analysis of the drought of 2019–2021. The analyses are based on the 1,800-year CESM preindustrial run with 94-, 98-, and 70-year segments, respectively (green), and the 35 Large Ensemble members (blue). The lengths of these segments correspond to the extent of the observed records used in CC2021, OK2021, and HW2022. The median values for the CESM preindustrial and Large Ensemble for each of the cases are indicated in the boxes. The median values are Australia: $\alpha_{CTRL}=4.9$, $\alpha_{LE}=0.4$; Siberia: $\alpha_{CTRL}=1.1$, $\alpha_{LE}=2.5$; Madagascar: $\alpha_{CTRL}=-0.02$, $\alpha_{LE}=0.09$.

### v3

@@ 2.5 Mean residual life plot
2.6 Hypothesis testing
Although WWA does not provide a detailed discussion of the hypothesis that they assume, we interpret that they adopt a null hypothesis of $\alpha$ equal to zero implying no ACC influence on extremes and an alternative hypothesis that, if $\alpha$ is not equal to zero, ACC does influence extremes. This inference is based on the quote, “If the one- or two-sided 95% confidence interval [for the change in probability, which has the same effect as $\alpha$] excludes zero, the change is statistically significant” (see Section 4.3.1 of Philip et al., 2020). We assess whether rejection of a null hypothesis of $\alpha$ equal to zero permits accepting an alternate of ACC influencing extreme events. In other words, we assess if a nonzero $\alpha$ necessarily means a role for ACC. We note that another null hypothesis for purposes of heat could be that the location of PDF of heat extremes is shifted by an amount equaling the GMST warming, as would be represented by $\alpha=1$. For purposes of consistency with the null implied by WWA, however, we only consider a null of $\alpha$ equal to zero.
WWA uses two variants of their analysis approach, one in which the extreme data in question is incorporated into the estimate of the PDFs, which we follow here, and one in which it is excluded Philip et al. (2020). The act of choosing to study a class of events immediately following the occurrence of a particularly extreme example is known to induce a bias related to whether it is included or excluded (Miralles and Davison, 2023, see). We choose to include the extreme value in question, which tends to make the parametric dependence on GMST larger and tends to make it easier to reject the null of no change in favor of the alternate hypothesis. Within the context of the WWA hypothesis test, this choice to include the extreme in question makes it more likely to conclude that AAC significantly influences extreme events.
@@ 2.7 Data
Following CC2021, we rely on two sets of observations: (1) daily maximum temperature extremes for each year from station data in Verkhoyansk as discussed in CC2021; and (2) 4-year smoothed GMST anomalies from GISTEMP Lenssen et al. (2019). CC2021 also performed a GEV analysis on temperature anomalies over a region in Siberia, which we do not focus on here in order to study specifically the extreme values detected at the Verkhoyansk station. At the time of the CC2021 study, the daily maximum temperature extreme in 2020 at the Verkhoyansk station was recorded as 38C. This temperature extreme currently appears as a missing data point on the National Climatic Data Center website (https://www.ncei.noaa.gov/cdo-web/search, station RSM00024266, VERHOJANSK, data downloaded Aug 2023) but we still use this unconfirmed value in our analysis.
We also study GMST anomalies and extreme temperatures over the Siberian region defined in CC2021 within the CESM Large Ensemble Kay et al. (2015) over the period 1926–2019. The CESM Large Ensemble contains 40 simulations of the climate over the period 1920–2100 under historical and RCP 8.5 external forcing, of which we use the 35 that were run on the NCAR supercomputer and not those runs done on the University of Toronto supercomputer that contain a global mean temperature bias National Center for Atmospheric Research (2024). These simulations differ only in small perturbations to the initial conditions, which lead to different temporal sequences of modes of internal variability. We use model output only up to the year 2020, and, as a result, there is little sensitivity to the emissions scenario used (DAMIP, Gillett et al., 2016, uses RCP 4.5) because emissions scenarios only significantly diverge later.
In addition, we study GMST anomalies and extreme temperatures within a 1,800-year CESM preindustrial simulation, yielding 18 segments of 94 years, chosen for consistency with the length of the observed record. This preindustrial simulation uses fixed greenhouse gas concentrations, volcanic activity, anthropogenic aerosols, land use, and solar forcing all set to their respective levels in the year 1850. Most importantly in the context of this study is the lack of time-varying anthropogenic forcing in this simulation.
Australian Bushfire Heatwave.
Following OK2021, we again use two sets of observations for the years 1920–2019: (1) annual (July–June) maxima of 7-day moving average daily-maximum surface temperature data from the Australian Water Availability Project (AWAP) over the Southeastern region in Australia defined in OK2021; and (2) 4-year smoothed GMST anomalies from GISTEMP. We apply a similar analysis to the CESM Large Ensemble and to the long CESM preindustrial run as for the Siberian Heatwave case except, in this case, using 18 segments of 98 years for consistency with the observations.
@@
Following HW2022, we use two datasets. The first data are 2-year means of precipitation from ERA5 Hersbach et al. (2020) over the region of Madagascar that was defined in HW2022. HW2022 explained that they used a “24-month running mean rainfall data from July to June.” Our approach is to calculate total annual precipitation and then calculate a running 2-year average of this annual data resulting in a new smoothed annual time series. For 1951–2020 the bottom 20% precipitation 2-year periods are selected for the analysis, corresponding to the following 14 two-year periods: 1956–1957, 1957–1958, 1958–1959, 1959–1960, 1962–1963, 1990–1991, 1991–1992, 1992–1993, 2008–2009, 2009–2010, 2015–2016, 2016–2017, 2017–2018, 2019–2020. The second data are 4-year smoothed GMST anomalies from GISTEMP. We note that the analysis following HW2022 relies on a fit to only 14 data points. We discuss the issues that are involved in fitting extreme distribution functions to such a small number of data points below. We also study the equivalent variables within the CESM Large Ensemble over the period 1950–2019, consistent with the period of observations. We again apply a similar analysis to the CESM Large Ensemble and to the long CESM preindustrial run, in this case using 18 segments of 70 years.
We note that in each of these three case studies, WWA leverages many datasets, including ERA5 reanalysis, station observations, and the multimodel CMIP5 ensemble. We use a subset of the data that WWA study but also the CESM Large Ensemble data in a consistent manner across the three example cases that we evaluate.
@@ 3 Results
We first describe our analyses of how extreme events depend on GMST in the context of the long CESM pre-industrial simulation in subsection 33.1. The simulation contains no time-varying anthropogenic forcing, making it useful for evaluating a null-hypotheses associated with no anthropogenic influence. We then consider each of the three case studies in subsections 33.2, 33.3, and 33.4. For each case we assess how GEV or GPD distributions fit the observations relative to normal or log-normal distributions. We then evaluate evidence for the inclusion of a GMST dependence within distributions using the deviance statistic and estimate the uncertainty level of the parameter $\alpha$ used for the attribution. Finally, we examine the amount of data needed to adequately constrain GMST dependence in a GEV or GPD fit.
3.1 Extremes analysis of a preindustrial simulation
@@
WWA examines the dependence of the distribution parameters of extreme events on the 4-year running mean GMST. Commonly a linear dependence of a location parameter on GMST, as represented by the parameter $\alpha$, is used to estimate the PDF of extreme events in, for example, 1900 versus 2020. Changes in the PDF between these epochs are interpreted as owing to the effects of anthropogenic climate change, implying that $\alpha$ indicates the extent to which anthropogenic climate change affects extremes. We show here that it is possible, however, for internal climate variability, such as ENSO, to produce covariance between GMST and extremes that is expressed as a non-zero $\alpha$ even without anthropogenic influence. This covariance is demonstrated most clearly in this section by using a pre-industrial climate model run, but we also elaborate upon this basic result using the CESM Large Ensemble. For attribution purposes this results suggests that the value of $\alpha$ calculated in the presence of a GHG increase can be affected by internal variability, biasing the attribution results. It should be noted that in a realistic scenario $\alpha$ is affected by both anthropogenic climate change and internal variability. The scenario of internal variability alone analyzed here is artificial, of course, and does not capture that, in the current climate, GMST is a reasonable proxy for anthropogenic climate change. Analyzing simulations that only include internal variability nevertheless provides insight into the effects of internal variability in more complete scenarios.
@@
To examine this potential role of internal variability, we start with the case of the hot weather associated with the bushfires in Southeastern Australia and perform a GEV analysis using a 1,800-year CESM preindustrial run. This run holds anthropogenic CO2, other anthropogenic greenhouse emissions, volcanic (natural) emissions, land use changes, and anthropogenic aerosols at values estimated for the year 1850. We divide the run into 18 data segments of length 94 years, equal to the length of the observed record and analogous to having 18 simulations of historical climate without ACC. The distribution of $\alpha$ values for the GEV preindustrial analysis is shown by the green bars in Fig. 1a. We find a large and positive $\alpha$ calculated from essentially all model output segments centered around a median value of 4.9 C per C, or that extreme summer temperatures increase by 4.9 C for every 1 C GMST warming. Note that, following WWA, we are using four-year-running average GMST values for this computation.
@@
It is difficult to imagine a physical causal mechanism that would lead to such an enormous amplification of the GMST signal in Australian heatwaves. A more plausible interpretation is that internal variability — for example, ENSO — could increase both extreme temperatures over Australia and, to a smaller degree, GMST, generating covariance between extreme events and GMST. In this scenario GMST is not an indicator of forcing of local extremes; rather, both local extremes and GMST are influenced by a third factor. It is known that several modes of internal variability affect extreme temperatures in Australia, including ENSO, the Indian Ocean Dipole, and the Southern Annular Mode Hendon et al. (2007); Ummenhofer et al. (2009); King et al. (2020). Min et al. (2013) find that ENSO and the Indian Ocean Dipole can lead to warming of extreme temperatures in Southeastern Australia by around $1^{\circ}$, indicating that internally generated variability can have an effect that is similar in magnitude to ACC.
WWA recognizes that $\alpha$ may be influenced by internal variability and, as noted, use a 4-year running mean GMST to filter out the effects of short-term internal variability. This approach does not fully remove the ENSO signal, however, because much of the variability is at longer periods. Spectral analysis of the annual-average Bivariate ENSO time series Smith and Sardeshmukh (2000) between 1870–2022, for example, indicates that 63% of the variance in the time series resides at frequencies below 1/(4 years). There is also the potential for nonlinear interactions between ENSO and GMST and the presence of slower modes of variability such as the Pacific Decadal Oscillation (PDO).
A similar analysis of Australian heat extremes made using the CESM Large Ensemble, which includes anthropogenic forcing, gives a distribution of $\alpha$ that is closer to zero (blue bars in Fig. 1) than in the unforced simulations (green bars). This indicates that both ACC and internal variability influence regional extremes and GMST, such that estimates of $\alpha$ will reflect multiple processes. In simulations with time-varying anthropogenic forcing, however, there is no way to disentangle the effect of internal variability on $\alpha$ using the WWA approach as described in published literature. We conclude that the magnitude of $\alpha$ is a biased indicator of anthropogenic influence. Including indices of internal variability as covariates may help to disentangle their respective contributions, though the projection of internal variability onto GMST can still confound the results.
These results highlight a potential major difficulty with analyses based on a fit of distribution parameters to GMST in a more realistic scenario that includes both internal variability and ACC: the calculated $\alpha$, and therefore the shift in the probability of extreme events, may be due to internal variability rather than only ACC. When analyzing the observational record in section 3.3, we find $\alpha\approx 2\;^{\circ}$ per C. This suggests a two-degree warming of extreme events for a one-degree warming of the GMST. Based on our results from the CESM preindustrial run with a time-invariant anthropogenic forcing, however, we suspect that the estimate obtained for $\alpha$ from the observed record is biased high by internal variability.
@@
Analysis of the CESM preindustrial run for the case of Siberian heatwaves (green bars in Fig. 1b) shows that the value for $\alpha$ deduced from a record similar in length to that of the observations ranges between 1 to 4, although there is again no time-varying anthropogenic forcing in the corresponding model run. This large and variable response implies that a non-zero $\alpha$ does not necessarily imply an effect of ACC. This analysis shows that $\alpha$ equal to zero is an unsuitable null hypothesis for the purposes of evaluating the presence of an anthropogenic influence on extreme events.
@@
For the case of Madagascar Drought (Fig. 1c) we find a large spread in $\alpha$ for both the preindustrial run as well as the CESM Large Ensemble, with both centered near zero. The fact that the $\alpha$ distribution from the preindustrial run is centered around zero suggests that a null hypothesis of $\alpha$ equal to zero may be appropriate, but the fact that $\alpha$ from the Large Ensemble is also centered around zero indicates that the there is little potential for this test to discern an ACC influence on extremes. That is, low extremes of precipitation show no systematic increase or decrease related to GMST in either the forced or unforced runs such that this test has low statistical power. We further consider this case below using observations.
@@
The analysis in this paper shows that internal variability can lead to a large $\alpha$. We conclude that in a more typical scenario analyzed using the WWA methodology and applied to observations under varying anthropogenic GHG and aerosols, the value of $\alpha$ may still be affected by internal variability and, therefore, may bias the estimates of the effect of ACC on the return time of extreme events. One could justify the WWA use of the 4-year averaged GMST as being a proxy for known greenhouse gas concentrations. If CO2 were indeed used, the analysis of the preindustrial run would, of course, show no dependence, eliminating this problem, though then introducing issues associated with lags between GHG concentrations and warming Proistosescu and Huybers (2017). Because WWA studies use GMST, we follow that methodology here.
@@
The result that internal variability may strongly affect $\alpha$ when the 4-year average GMST is used to link extreme events to ACC suggests that it might be useful to reevaluate some of the WWA results based on the fit to extreme distributions in at least the three studies that we review. It is possible, of course, that a strong anthropogenic signal in a few decades could significantly affect extreme events and overwhelm the effects of internal variability in cases such as the Siberian Heatwave. The attribution of extreme temperatures may be more difficult in cases like the Australian example, where the effects of internal variability are large, as indicated by the large $\alpha$ detected in the CESM preindustrial simulation. In such cases, it would be important to correctly identify all relevant covariates — such as the possible effect of ENSO in the Australian case — and note that misleading results for $\alpha$ may be obtained if one relies only on GMST.
@@
Figure 1: Effect of internal climate variability on $\alpha$ distribution. PDFs of $\alpha$ for (a) The Australian 7-day moving mean of the maximum temperature related to the analysis of the 2019–2020 bushfires. (b) The Siberian maximum daily temperatures that were used in the analysis of the 2020 heatwave. (c) The Madagascar 2-year precipitation used in the analysis of the drought of 2019–2021. The analyses are based on the 1,800-year CESM preindustrial run with 94-, 98-, and 70-year segments, respectively (green), and the 35 Large Ensemble members (blue). The lengths of these segments correspond to the extent of the observed records used in CC2021, OK2021, and HW2022. The median values for the CESM preindustrial and Large Ensemble for each of the cases are indicated in the boxes. The median values are Australia: $\alpha_{CTRL}=4.9$, $\alpha_{LE}=0.4$; Siberia: $\alpha_{CTRL}=1.1$, $\alpha_{LE}=2.5$; Madagascar: $\alpha_{CTRL}=-0.02$, $\alpha_{LE}=0.09$.